OpenStudy (anonymous):

simplify 5x / x^2+4x

OpenStudy (anonymous):

is your question written like this \[\huge \frac {5x}{ x^2+4x}\]

OpenStudy (anonymous):

\[\huge \frac {5x}{ x^2+4x} =\huge \frac {5x}{ x(x+4)}=\huge \frac {5}{(x+4)}\] @pilla